40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 52 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  231   Wed Aug 5 16:59:45 2009 DmassComputingDAQTransfer Functions and DTT

Quote:

A working servo loop should not have such a low openloop gain like -60dB or -80dB in the working frequency band.
Please check the setting and the math.

Ultimately any measured openloop tf should be explained by the interferometer model, the filter response, and the actuator response.
(e.g. Ph.D thesis by M.Ando P.142 http://t-munu.phys.s.u-tokyo.ac.jp/theses/ando_d.pdf )

Quote


Quote:

Oh...This should not be the openloop transfer function. You must multiply the transfer function between two measurement points
unless those two points are just before and after the injection point.

 The channels are set up in such a way that I am measuring on either side of my injection  point with no filters in between

 

 

 

Now that we have DTT fully up and running with AWG, I took a *good* open loop transfer function for the PMC control loop.

I am sure there are some things I haven't yet understood about my loop...

In the digital world I have:

  • A 1 Hz pole for control
  • Two 1Hz zeros and 10 Hz poles for antidewhitening to comensate for the Piezo Driver Box.
Attachment 1: PMCOLTFSWEEP.jpg
PMCOLTFSWEEP.jpg
  230   Wed Aug 5 15:14:55 2009 DmassComputingCDSAWG and TP now up

I emailed Alex and:

 

Hi David,
looks like you have switched from running om1/om2 pair to running atf.
When you do that you have to alter testpoint.par file:


[controls@oms param]$ pwd
/cvs/cds/caltech/target/gds/

param
[controls@oms param]$ ls -alt testpoint.par
-rw-r--r-- 1 controls controls 330 Aug  5 08:52 testpoint.par

I changed om1 to atf there.

Testpoints are working now.
 
Test points and AWG are now up.

  229   Wed Aug 5 10:10:11 2009 MichelleMiscGeneralWeekly Update (7?)

A number of things made the past week somewhat unproductive, until yesterday. I was at a conference Mon-Wed. last week and was sick that Friday. This Monday I finished my second progress report (now on the wiki).

Yesterday we made major progress in the setup and alignment of the cavity. It was really helpful to have optics people (Zach and Frank) down in the lab, I learned a lot about alignment. We replaced the two lenses we had initially (one before the FI and one before the EOM) with one slowly converging lens which would place the waist in the EOM. Frank also said there was no need to have two steering mirrors before different optics, since our laser height is fixed. We were initially going to use a curved mirror for the high reflectance mirror at one end of the cavity, but Rana suggested that we would get a lock faster if we just used two flat mirrors with a lens in between them (since the finessed of the cavity doesn't really matter right now). We're also not going to worry about mode-matching just yet.

The cavity is now roughly aligned. It will need some fine-tuning, but we should be able to start trying to lock today. We're going to start with an analog system to get a feel for the parameters and just to get some sort of error signal, and then later we'll switch everything over to the DAQ. That's about all for now.

  228   Tue Aug 4 23:52:06 2009 KojiComputingDAQTransfer Functions and DTT

A working servo loop should not have such a low openloop gain like -60dB or -80dB in the working frequency band.
Please check the setting and the math.

Ultimately any measured openloop tf should be explained by the interferometer model, the filter response, and the actuator response.
(e.g. Ph.D thesis by M.Ando P.142 http://t-munu.phys.s.u-tokyo.ac.jp/theses/ando_d.pdf )

Quote:

Quote:

Oh...This should not be the openloop transfer function. You must multiply the transfer function between two measurement points
unless those two points are just before and after the injection point.

 The channels are set up in such a way that I am measuring on either side of my injection  point with no filters in between

 

Attachment 1: Pages_from_ando_d.png
Pages_from_ando_d.png
  227   Tue Aug 4 21:55:00 2009 DmassComputingDAQTransfer Functions and DTT

Quote:

Oh...This should not be the openloop transfer function. You must multiply the transfer function between two measurement points
unless those two points are just before and after the injection point.

Quote:

Quote:

So I was trying to take some transfer functions with DTT using an analog sweep.

We don't have AWG or test points, so I was trying to do it with DAQ channels. I set number of A channels as 1, set the rest, then tried to run it.

I get the error "Readback channel unavailable (); Unable to setup test channels" 

I may be SOL to take any transfer functions without either AWG or the test points working if there is some sort of handshake missing.

 

Using the Fourier tools, I took a first attempt at getting a good open loop transfer function for my PMC.

The bottom right plot is my error signal in blue, and my noise injection in red.

 

 The channels are set up in such a way that I am measuring on either side of my injection  point with no filters in between

  226   Tue Aug 4 19:33:47 2009 KojiComputingDAQTransfer Functions and DTT

Oh...This should not be the openloop transfer function. You must multiply the transfer function between two measurement points
unless those two points are just before and after the injection point.

Quote:

Quote:

So I was trying to take some transfer functions with DTT using an analog sweep.

We don't have AWG or test points, so I was trying to do it with DAQ channels. I set number of A channels as 1, set the rest, then tried to run it.

I get the error "Readback channel unavailable (); Unable to setup test channels" 

I may be SOL to take any transfer functions without either AWG or the test points working if there is some sort of handshake missing.

 

Using the Fourier tools, I took a first attempt at getting a good open loop transfer function for my PMC.

The bottom right plot is my error signal in blue, and my noise injection in red.

 

  225   Tue Aug 4 18:41:49 2009 DmassComputingDAQTransfer Functions and DTT

Quote:

So I was trying to take some transfer functions with DTT using an analog sweep.

We don't have AWG or test points, so I was trying to do it with DAQ channels. I set number of A channels as 1, set the rest, then tried to run it.

I get the error "Readback channel unavailable (); Unable to setup test channels" 

I may be SOL to take any transfer functions without either AWG or the test points working if there is some sort of handshake missing.

 

Using the Fourier tools, I took a first attempt at getting a good open loop transfer function for my PMC.

The bottom right plot is my error signal in blue, and my noise injection in red.

Attachment 1: OMCOLTF2.jpg
OMCOLTF2.jpg
  224   Tue Aug 4 15:46:46 2009 DmassLab InfrastructureHVACTemperature update

It has been a brisk 72 degrees for a while (~1 week?) in lab. This is apparently the proper temperature for keeping graduate students awake, and lasers cool.

  223   Tue Aug 4 15:37:17 2009 DmassComputingCDSNokia Wireless Device

Quote:

To use the small Nokia wireless device:

 

1) Connect to cdsrana

2) Check what the ip of the machine you want to use is (/sbin/ifconfig)   - 131.215.113.56 at the time of this entry

3) run an x-terminal on the nokia (under extras)

4) enter into the command line :   ssh controls@IP

5) It will prompt you for a password, enter.

you can now run simple medm screens of your design from anywhere with wireless access to the cdsrana network!

 

The Nokia wireless device was not able to connect to ws1, and for some reason, (probably related ?) we couldn't ssh into ws1 from the outside world either.

I reset the routers, and renewed the DHCP lease. This seems to have fixed the problems.

New IPs:

131.215.113.56 (for cds - same)

131.215.114.120 (for the outside world)


I have also sent Alex an email asking him to fix our awg/test points now that he can ssh in.

  222   Tue Aug 4 15:01:28 2009 DmassComputingDAQTransfer Functions and DTT

So I was trying to take some transfer functions with DTT using an analog sweep.

We don't have AWG or test points, so I was trying to do it with DAQ channels. I set number of A channels as 1, set the rest, then tried to run it.

I get the error "Readback channel unavailable (); Unable to setup test channels" 

I may be SOL to take any transfer functions without either AWG or the test points working if there is some sort of handshake missing.

  221   Mon Aug 3 22:53:11 2009 DmassLaserPMCPhotodiodes and Power meters - confusion

I am measuring the power of my PMC transmission with a Thorlabs PDA10CS. I have one after a 50:50 Beamsplitter on the output.

 

I measured 603 mV through a 2.0 ND filter on a 0dB gain setting with an oscilloscope (1MOhm load). This should be half my transmitted power.

 

I calculate: 0.603 V *(1A/1.5*10^3V)*(1W/.7A)*100 (for ND filter) = 58 mW...

 

I measure again with a 10 dB setting for sanity:

1.94 V * (1/4.75*10^3)/\*(1/.7)*100 = 58 mW....

 

BUT when I put a power meter in the same part of the beampath, I get a reading of 30 mW. My input power is 90 mW...I think I am getting at least 2/3 transmission, but am confused by my photodiode reading. It seems to imply I am getting more power out than I put in. (hmm.)

  • ARE THERE ANY TRICKS TO GETTING THE POWER METER TO BEHAVE?
  • OR: IS THERE SOMETHING STUPID I AM DOING IN MY CALCULATION
  220   Mon Aug 3 22:41:26 2009 DmassElectronicsPMCBlood Feud Over

Quote:

The war is now at an end. Aidan is now free to use the Marconi, and I am using the SRS DS345 we picked up to generate 13 dBm of 30 MHz to lock with.

In other news, Koji & I discovered that there were two dewhitening filters in the piezo driver box which were not being compensated. I put in anti dewhitening (1:10 and 1:10) at the last filter bank before the DAC, and a 1 Hz pole (:1). I was then able to lock with high gain (see attached spectrum). Koji says "it looks like your UGF is around 400 Hz" based on the spectrum. I have included a PDF for more detail.

The Red, Blue, and Green are the plots of PMC Transmission with different loop gains. I chose green (14000) to be my operating point.

When I swapped out the Marconi for the DS345 I took another spectrum of transmission with 14000 gain. That is the brown curve. I am now happy using the DS345*

 

 

*subject to change without notice

 

Spectrum is PMC ERROR SIGNAL WITH VARIOUS LOOP GAINS

 

 

 

I also took some spectra of the tranmission with a new alignment and a new lock.

Attachment 1: PMCTransGainsreal.jpg
PMCTransGainsreal.jpg
  219   Mon Aug 3 21:23:46 2009 Aidan, ConnorLaserFiberSuccess! Fiber noise suppression ... again (but not calibrated)

We added an out-of-loop PD (New Focus 1811) on Friday and measured the added phase noise of the fiber. This allowed us to take a demodulated output from one PD and feed it back to the VCO (Marconi) to actuate on the frequency driving the AOM and use the new PD as witness measurement of the noise reduction.

The attached spectra show a reduction in the measured phase noise (well, it's not really phase noise until it's calibrated) for a variety of different FM Deviation (transductance) settings (Hz/Volt) on the Marconi. The largest value we could get, 800kHz (per Volt?), gave us the most suppression of the fiber noise.

I'll attach a layout of the electronics, including all the gain settings on SR560s, in the next day or so.

The next thing we need to do is to make sure that all the cables we have carrying the RF signals are the right lengths to give the correct phases at each mixer and at the AOM. Simultaneously, we need to calibrate everything very carefully now.

(I'm curious as to why the signal was much larger at an FM Devn of 100kHz. Also, I'm not sure why the 800kHz spectra saturates at a lower level than the free-running noise. I checked the DC output of the PD with an without the feedback and it appeared to be unchanged.

Morale of the story: we need to fully understand the system we have before we move any further forward.

Attachment 1: 2009-08-03_noise_reduction.pdf
2009-08-03_noise_reduction.pdf
  218   Mon Aug 3 17:32:52 2009 DmassElectronicsPMCBlood Feud Over

The war is now at an end. Aidan is now free to use the Marconi, and I am using the SRS DS345 we picked up to generate 13 dBm of 30 MHz to lock with.

In other news, Koji & I discovered that there were two dewhitening filters in the piezo driver box which were not being compensated. I put in anti dewhitening (1:10 and 1:10) at the last filter bank before the DAC, and a 1 Hz pole (:1). I was then able to lock with high gain (see attached spectrum). Koji says "it looks like your UGF is around 400 Hz" based on the spectrum. I have included a PDF for more detail.

The Red, Blue, and Green are the plots of PMC Transmission with different loop gains. I chose green (14000) to be my operating point.

When I swapped out the Marconi for the DS345 I took another spectrum of transmission with 14000 gain. That is the brown curve. I am now happy using the DS345*

 

 

*subject to change without notice

 

Spectrum is PMC ERROR SIGNAL WITH VARIOUS LOOP GAINS

Attachment 1: PMCTRANSLoopGains.pdf
PMCTRANSLoopGains.pdf
Attachment 2: PMCTransLoopgains.jpg
PMCTransLoopgains.jpg
  217   Sun Aug 2 22:48:53 2009 DmassLaserPSLPMC Locked - 10% Transmission

I grabbed some lenses from the nether regions of the lab, and have managed to get the PMC relocked. I did a cursory alignment, and 10% transmission seems to be my max for now. I haven't optimized my mode matching because I am changing lenses tomorrow. I spent most of the night trying to optimize my alignment and getting a good number for transmitted power.

 

 

I was still:

  • Saturating my photodiode on the PMC reflection by not dumping most of the beam. I put a 90% splitter in front of it w/ a dump
  • I am sure there are more things.
  216   Sun Aug 2 22:43:44 2009 DmassLaserPSLPMC Locking woes

Quote:

So it turns out that I was putting my 35MHz signal from the Marconi (13 dBm before a spilitter) without any impedance matching and probably causing all kinds of havoc with my RF reflections.

I changed the setup so I have a 50ohm resistor T'ed into the input for the EOM at the PSL enclosure. My error signal is now 10 mV rather than 1ish without the 560.

*** check RF phase ***

 Each time I changed something I put in 90 degrees of phase via cable into my mixer to make sure I haven't unintentionally fubared the phase

  215   Sun Aug 2 22:41:23 2009 DmassElectronicsGeneralPhotodiode Woes

The PDA10CS from Thorlabs appears to have a peculiarity, or I am just confused. The Spec Sheet says that it should have a 10V max output for a Hi Z load. This appears to not be true for its 0 dB gain setting. It rails at 5.44V with zero dB of gain, and 10.something V at higher gain settings. This caused some modicum of confusion.

 

It was plugged into a Tektronix 1001B which is high Z (I measured to double check - 1MOhm).

 

Is it fine to use ND filters with my photodiodes? It seems fine, but I don't know if there is somet hing I am missing there.

  214   Sun Aug 2 17:01:45 2009 DmassLaserPSLPMC Locking woes

So it turns out that I was putting my 35MHz signal from the Marconi (13 dBm before a spilitter) without any impedance matching and probably causing all kinds of havoc with my RF reflections.

I changed the setup so I have a 50ohm resistor T'ed into the input for the EOM at the PSL enclosure. My error signal is now 10 mV rather than 1ish without the 560.

*** check RF phase ***

  213   Fri Jul 31 14:58:14 2009 KojiLaserPSLReReReLocking the PMC aaaaaand stuck.

We inserted an SR560 with G=+100 as a preamplifier for the mixer output so that we can have a ~100mV error signal for a cavity swing.
The error signal of ~mV with the offset of ~mV (presumably coming from both the ADC and the mixer) is a totally wrong approach.

Ah, you don't need the 50Ohm termination at the ADC input any more as we are handling only AF signals there.
It just make the signal amplitude half.

Quote:

I called upon Koji to assist me in my woes.

We got a lock by turning off the digital filter. We could not lock with it on at any gain we tried

I will take a transfer function with the 35670A and try to get at the real transfer function. 

Sure would be nice to have AWG

 

  212   Fri Jul 31 14:38:51 2009 DmassLaserPSLReReReLocking the PMC aaaaaand stuck.

Quote:

I spent some time tonight realigning and relocking (attempting to relock) the PMC.

I got to a point where I did not understand the behavior of my error signal. I did the standard things...

  • Changed cable length by pi/2
  • Eventually realized that I had done this all before and looked at my old elog to not do the same dumb things again
  • Same setup as this elog entry
  • Change the offsets muchly
  • Change the gain
  • Change the sign of the gain
  • Make sure just a single pole @ 1Hz is toggled on

I get a behavior where my transmitted beam would oscillate in power, seemingly about the center, and my error signal looked as below

 

When I say error signal I mean "Thing that should be my error signal but for some reason is not" this is the raw signal I am inputting into the DAQ - so this is what I pass to my filters then to teh PZTChange the offsets

 

I called upon Koji to assist me in my woes.

We got a lock by turning off the digital filter. We could not lock with it on at any gain we tried

I will take a transfer function with the 35670A and try to get at the real transfer function.

 

Sure would be nice to have AWG

  211   Fri Jul 31 12:27:19 2009 Aidan, ConnorLaserFiberAdjusting angle of fiber-input-HWP to minimize fluctuations in the output

The attached time series shows measurements of the power in the s and p polarizations (via ThorLabs PDA10CS PDs) for different input HWP settings. This just shows the raw data from the experiment  described here ...

Attachment 1: 2009-07-31_matching_input_polarization.pdf
2009-07-31_matching_input_polarization.pdf
  210   Fri Jul 31 01:12:41 2009 DmassLaserPSLReReReLocking the PMC aaaaaand stuck.

I spent some time tonight realigning and relocking (attempting to relock) the PMC.

 

I got to a point where I did not understand the behavior of my error signal. I did the standard things...

  • Changed cable length by pi/2
  • Eventually realized that I had done this all before and looked at my old elog to not do the same dumb things again
  • Same setup as this elog entry
  • Change the offsets muchly
  • Change the gain
  • Change the sign of the gain
  • Make sure just a single pole @ 1Hz is toggled on

I get a behavior where my transmitted beam would oscillate in power, seemingly about the center, and my error signal looked as below

 

When I say error signal I mean "Thing that should be my error signal but for some reason is not" this is the raw signal I am inputting into the DAQ - so this is what I pass to my filters then to teh PZTChange the offsets

Attachment 1: ErrSignal.png
ErrSignal.png
  209   Thu Jul 30 21:39:49 2009 Aidan, ConnorLaserFiberFIber Input HWP set to 356.4 degrees.

Aidan and I matched the polarization of the beam to the polarization-maintaining fiber.

To begin with, we sent the output beam through a polarizing beamsplitter and directed both the transmitted (P-polarized) and reflected (S-polarized) light into photodetectors. We oriented the output fiber axes such that the reflected (S-polarized) power is minimized.

Next, we wanted to orient a HWP at the input of the fiber such that fluctuations in the reflected (S-polarized) output beam are minimized. This would mean that the polarization of the incoming light is matched closely to the fast axis of the fiber, because fiber length fluctuations don't cause polarization changes in the output beam.

To find the amplitude of fluctuations in the reflected (S-polarized) output, we aimed a heat gun onto the fiber, lengthening it by multiple wavelengths so that the photodetector output reached clear minimum and maximum values. We recorded the difference in these values.  I will soon attach a plot of the reflected output amplitude fluctuations vs HWP angle.

Attached are photographs of the current input and output layouts.  The angle with minimum output fluctuations is 356.4 degrees.

Attachment 1: Fibre_Polarization_Matching_Input.jpg
Fibre_Polarization_Matching_Input.jpg
Attachment 2: Fibre_Polarization_Matching_Output.jpg
Fibre_Polarization_Matching_Output.jpg
  208   Wed Jul 29 19:59:15 2009 Aidan, ConnorLaserFiberHWP inserted to match input polarization to fast axis of PM fiber

We put a HWP immediately before the input to the fiber to match the polarization to the fast axis of the fiber. We can see the effect of this in the contrast of the fringes in the demodulated AC output of the interferometer. As the fiber expands and contracts the output polarization becomes more or less circularly polarized depending on how much the input polarization is mis-aligned from the fast axis. Below is a time series of the demodulated output of the PD over about 45 minutes. We need to optimize the position of the HWP but that's a job for tomorrow.

 

Attachment 1: 2009-07-29_wandering_polarization.pdf
2009-07-29_wandering_polarization.pdf
  207   Wed Jul 29 18:04:54 2009 Aidan, ConnorLaserFiberNoise Budget for Single-Pass Setup

Aidan and I generated power spectra of uncalibrated phase noise, reference only intensity noise, fiber only intensity noise, DAQ noise, and dark noise using the DTT to create a noise budget.  We made a shot noise calculation and imported it onto the same plot.

The intensity noise of each arm seems to be limited by dark noise. The spectra are attached.

Our next goal is to match the polarization of the fiber arm beam to the fiber's fast axis.

Attachment 1: 2009-07-29_PD1_Demodulated_output_noise_budget.pdf
2009-07-29_PD1_Demodulated_output_noise_budget.pdf
  205   Wed Jul 29 16:57:26 2009 Aidan, ConnorLaserFiberCurrent fiber electronic setup

 The current demodulation setup:

Attachment 1: FS_layout_2009_07_29.png
FS_layout_2009_07_29.png
  204   Wed Jul 29 16:39:00 2009 Aidan, ConnorLaserFiberInterference signal setup for FS

Aidan and I are acquiring a beat signal from the interference of a reference beam and a frequency-modulated beam. The demodulated output can be seen on THIS ENTRY.

I have attached a photograph of our setup below. The reference beam is a pickoff from the 1W NPRO, and the zeroth order AOM beam is blocked by an iris.

Aidan will shortly post a schematic of the electronics, and we hope to start actuating on the laser frequency later today.

Attachment 1: Single_Pass_Noise_Measurement_Setup.jpg
Single_Pass_Noise_Measurement_Setup.jpg
  203   Wed Jul 29 15:46:38 2009 Aidan, ConnorLaserFiberOutput from demodulated AC channel in FS expt - polarization not totally matched

 

Connor and I have set up an 80MHz optical beat on the photodetector in the fiber setup, (one beam is passed through an AOM to freq. shift, is passed through the fiber and then recombined with a reference), demodulated the output from the PD and acquired it in the DAQ.

A 3 hour+ time series of the demodulated output is shown below. The variation in the maxima and minima over the 3 hours is most likely due to fluctuations in the output polarization state from the fiber due to the input not being completely matched to th fast axis of the PM fiber.

We're going to insert a HWP before the fiber to adjust the input polarization so that it is aligned with the axis of the fiber.

Attachment 1: 2009-07-29_PD1_Demodulated_output.pdf
2009-07-29_PD1_Demodulated_output.pdf
  202   Tue Jul 28 18:30:29 2009 ConnorLaserFiberFS Reference Beam Calibration

Aidan and I wanted to know the Mach-Zehnder reference beam power as a function of the angle of a half-waveplate which, coupled with a polarizing beamsplitter, attenuates the beam. Between the half-waveplate and the Mach-Zehnder arm, the reference beam is attenuated by a 50/50 beamsplitter.

Here is our data:

Angles (degrees from vertical of fast axis): [30, 36, 42, 48, 54, 56, 58, 60, 62, 64, 66, 68, 70, 72, 74, 76, 80, 86, 92, 98, 104, 110]

Power (mW): [397, 328, 237, 146, 70, 47, 31, 17, 7, 4, 2, 7, 14, 26, 41, 60, 110, 197, 291, 372, 430, 459]

I've attached a plot of our data with a least squares fit of a cos^2 function parametrized by amplitude, frequency, phase shift, and offset.

Attachment 1: Ref_Power_vs_HWP_Angle.pdf
Ref_Power_vs_HWP_Angle.pdf
  201   Mon Jul 27 17:57:55 2009 DmassLaserPSLMore Mode Matching

Quote:

Our lens positions for mode matching into the PMC seem to be based ON LIES. I cannot locate my old data on the first PSL scans and fitting. I have now learned a very important lesson in imitating Aidan.

Since my old "measurement" (quoted number) seems to disagree with what I have found by roughly a sign, I am operating under the assumption that I was simply wrong in the linked entry.

I have plotted our old mode matching solution with the new beam parameters.

I have also found 2 mode matching solutions for the "average" mode coming out of the PSL, one with two 500mm lenses, and an alternative in case we don't have them around.

 

If these do not suffice for our locking needs, we will need to deal with the ellipticity of the beam. I think it will be fine for now based on this.

 

I have included plots of the two offered mode matching solutions, as well as the picture of what we were using for coupling into the cavity.

 I checked the relative stability of each of the two mode matching solutions here by perturbing the lens positions and focal lengths.

 I used the Anderson formulas <add link> to calculate the power lost to first order into higher order modes. I then altered parameters (lens positions and focal lengths) as noted in the following tables and compared changes in power loss.

 

Solution with f1 = 0.5m   x1 = 0.42m     f2 = 0.5m    x2 = 0.62m    with positions relative to PSL enclosure
Change in Power Loss dx1 = 5mm dx2 = 5mm df1 = 1% df2 = 1% Absolute Loss
Red (Horizontal)  1.9%  0.29%  0.34%  1.4% 10.7%
Green (Average)  0.02%  0.02%  ~0%  ~0% 0.01%
Blue (Vertical)  0.44% 0.09%  0.3% 0.38% 3.38%
Solution with f1 = 1.0m   x1 = 0.47m     f2 = 0.4m    x2 = 0.55m    with positions relative to PSL enclosure
Red (Horizontal) 1.15% 0.76% 0.43% 1.4% 10.07%
Green (Average) 0.02% 0.02% 0.01% 0.06% 0.01%
Blue (Vertical) 0.31% 0.09% 0.14% 0.46% 3.67%

 

 

 

 

 

 

 

 

I choose the .5m and .5m solution.

 For lenses, I plan on using a plano-convex lens for as lens 1, and a biconvex lens as lens2.

 

  200   Fri Jul 24 19:00:37 2009 ranaLaser I've done some things in lab.
Make sure to use only ~5 mW to do all of the alignment.

Turn off the room lights if the power meter is not sitting well at zero.

Instead of the IR viewer, you can use a CCD camera to look at the input to the EOM and its output. You want to just center the beam in the EOM's apertures and then we will do the fine alignment by monitoring the RF AM.

Don't run the cavity with high power yet. Just 5 mW is easily high enough.
  199   Fri Jul 24 16:59:10 2009 ConnorLaserFiber1W NPRO minimum extinction ratio

The 1W NPRO beam is elliptically polarized. We use a quarter-waveplate (QWP) to linearize the polarization.

7 days ago, I made some measurements of the extinction ratio of the beam for various orientations of the QWP (attenuating the beam with a half-waveplate and polarizing beamsplitter) to find the QWP orientation at which the light is linearly polarized. The minimum ratio occured when the fast axis was between 10 and 15 degrees from the vertical. I've attached a plot of the data, and I will fit it to a curve as soon as I work out which function should describe the data.

Attachment 1: ExtRatio1WNPRO.pdf
ExtRatio1WNPRO.pdf
  198   Thu Jul 23 16:55:24 2009 Aidan, ConnorLaserFiberFiber coupling dominance - 19% transmission.

10:10AM : need to align AOM and check RF drive into it.
    AOM - wants < 2W RF Power Level
    Amplifier gain [ZHL-1A http://www.minicircuits.com/pdfs/ZHL-1A.pdf] - 16dB
    Maximum power output (dBm) = +28dBm = 630mW

14:00: Have AOM aligned. 3Vpp amplitude 80MHz signal to power amp gives ~63% diffraction efficiency [-1 order: 100+/-3mW, total: 160+/-3mW]

16:40: -1 order is aligned into fiber - coupler is very touchy and sensitive to misalignments. looks to be good transmission though. 37mW incident, 7mW transmitted - about 19%!

  197   Thu Jul 23 16:31:28 2009 MichelleLaser I've done some things in lab.

Update:

Today I aligned the laser beam through the EOM with something resembling a normal shape at the output. This was tricky. The powermeter isn't giving me very reasonable readouts, the aperture on the EOM is impossible to see with the IR viewer, the EOM is in the middle of the table where I can't really reach it, and it's very close to a focusing lens. The point being that I may have to tweak it a bit, but I really don't want to have to redo the alignment of that particular optic. So no earthquakes for a while.

A note on the power meter: It's been registering anywhere between -6 mW and 11 mW with no beam on it at all. I've been zeroing it before putting it in the beam before each use, but I don't know how much I trust it.

The beam shape coming out of the EOM  is still a bit funny - with one bright beam, and then a very faint ring, so it looks like a ring you would wear on your finger with a diamond on top. I am attributing this to the fact that the beam is slightly larger than the EOM's aperture; I have set it up so the center of the EOM coincides with the beam's waist. I'm also (according to the power meter) getting a high attenuation inside the EOM, but whether this is correct or normal I haven't yet figured out.

We need another mount for a beam splitter, this will go at the input to the cavity. We are also still in need of a quarter waveplate, and Aidan says he has a spare that he won't need to use for a few weeks. Also, CVI doesn't specify the damage threshold for the cavity mirrors for cw lasers, only for pulsed sources. Alastair has sent them an email requesting this information, and once we know that we can send the beam into the cavity at appropriate power levels.

  196   Thu Jul 23 16:05:48 2009 DmassLaserPSLMore Mode Matching

Our lens positions for mode matching into the PMC seem to be based ON LIES. I cannot locate my old data on the first PSL scans and fitting. I have now learned a very important lesson in imitating Aidan.

Since my old "measurement" (quoted number) seems to disagree with what I have found by roughly a sign, I am operating under the assumption that I was simply wrong in the linked entry.

I have plotted our old mode matching solution with the new beam parameters.

I have also found 2 mode matching solutions for the "average" mode coming out of the PSL, one with two 500mm lenses, and an alternative in case we don't have them around.

 

If these do not suffice for our locking needs, we will need to deal with the ellipticity of the beam. I think it will be fine for now based on this.

 

I have included plots of the two offered mode matching solutions, as well as the picture of what we were using for coupling into the cavity.

Attachment 1: OldMMSol.pdf
OldMMSol.pdf
Attachment 2: MMSol500500.pdf
MMSol500500.pdf
Attachment 3: MMSol1000400.pdf
MMSol1000400.pdf
  195   Thu Jul 23 13:14:03 2009 Aidan, ConnorLaserFiberAOM Aligned, beam matched into fiber, pick off from beam splitter correct size

 

Connor and I have been assembling the FS experiment. The following diagram shows the layout of everything that is aligned at the moment.

We've got the AOM installed and have it aligned to give -1 order when driven at 80MHz (attached image shows main 0th order and -1 order to the right. Not sure what spot up the top is).

Attachment 1: FS_layout_2009_07_23.png
FS_layout_2009_07_23.png
Attachment 2: photo.jpg
photo.jpg
  194   Thu Jul 23 07:33:41 2009 AidanComputingCDSAdded C2 MEDM screens to 40m SVN.

 

I've added our medm screens from the ATF to the 40m SVN. They can be found in 

https://nodus.ligo.caltech.edu:30889/svn/trunk/medm/c2/atf/

  193   Thu Jul 23 07:22:35 2009 AidanComputingCDSChecked out the 40m MEDM screens from SVN to /cvs/cds/caltech/medm/c1

 

As title says ...

  192   Wed Jul 22 20:40:13 2009 AidanLab InfrastructurefubarFiling cabinet lock engaged ... with no key

This happened mysteriously and had absolutely nothing to with me. The fact that I was the last person to open the filing cabinet before this happened is circumstantial and beside the point.

Will get the lockshop onto this in the next couple of days. In the meantime, just try and exercise your clairvoyance.

  191   Wed Jul 22 20:36:37 2009 Aidan, ConnorLaserFiberAligned beam into fiber ... mode-matching needs a little work

Connor and I got the beam aligned into the fiber today. After a little optimization of the fiber coupled XYZ I'd estimate we were getting maybe 10-15% transmission. I'm pretty sure we can improve this with some work. We caculated the input mode (which I will reproduce elsewhere) but it's looking for a mode size of approximately 6.6 microns diameter. With a 10X microscope objective (f = 16.5mm) we want the intput beam to the coupler to be around 3.4mm diameter - which we have.

We mounted the AOM and inserted it into the setup. We should be able to get an 80MHz frequency shifted beam through the fiber tomorrow measure the fiber noise. That same setup can be used in a PLL to suppress the fiber noise.

 

  190   Wed Jul 22 11:09:45 2009 MichelleMisc Weekly Update 5

Over the course of the past week I've done a few things. When I began alignment I discovered that our laser was periodically shutting itself down. This was a very perplexing problem for about 2 days before Alberto came in with the diagnosis: the diode was overheating. We'll be sending that in to get it fixed, and we'll also put a heat sink on its casing. Hopefully that will be up and running by the time we really need it.

Right now, Aidan's laser is set up with a 50/50 beamsplitter, and then a half wave plate at each output of that. That way we can run the laser at full power, and each of us can independently adjust the power going into our respective experiments. This is working well while we're aligning things, but it clearly won't work long-term - we need to act directly on the laser's frequency to lock it to a cavity. This may not fare so well for fiber noise suppression.

We have our setup mostly aligned. The beam is going through some steering mirrors, through a lens, into the Faraday Isolator, through another set of steering mirros and a lens, and into the EOM. There is very little loss inside the isolator (putting in ~35 mW and getting out 33-34 mW), however I have not been able to get the power at the output of the EOM higher than ~ 24 mW. I don't think this is normal, but I will check that with people who know better than I do. I think it is probably the fault of poor alignment - the aperture is ~ 2mm and it's about 2 cm from the lens in the middle of the table, so it's hard to reach it or even view it properly to see what needs tweaking.

Over the next week I plan to finish the alignment and hopefully get a lock. I'll get a picture up once the rest of the setup is in place and aligned.

 

Oh, and I also helped clean the lab this week. It's pretty shiny, except for the heaps of garbage boxes now sitting in the hallway. We'll take care of that soon.

  189   Wed Jul 22 09:43:53 2009 ConnorMisc Weekly Report #5

Most of this week has been spent aligning optics for the fiber stabilization experiment.  Part of the setup is shown here

The next step is to couple the beam into the fiber, single pass it and interferometrically beat it with the original signal so we can characterize the acquired phase noise.

Mode matching to various optical elements involves changing the beam parameter (i.e., beam width and radius of curvature) to fit specifications. We do this using lenses. Given the incoming beam parameter, the desired outgoing beam parameter, and lens focal lengths, we can find solutions (lens positions) to the problem.

  188   Tue Jul 21 20:41:56 2009 DmassLaserPSLRazor Blades!

Quote:
Yeah, real micrometer with a reading is > plain screw.


Converting from flemish ells was hard. I'll just use a mic next time.
  187   Tue Jul 21 20:39:18 2009 DmassLaserPSLPSL Beam Scans

I took a bunch of scans of the PSL with the Beamscan and compared them with some razorblade occlusion measurements

The first steering mirror after the beam exits the PSL was removed, and data was taken at this point in the beam path.

The numbers for the waist derived from my razorblade measurements are on the plot in green. The disagreement seems not horrible to me.

 

We should still fix the beamscan, and think about buying another (newer) one.

 

I am told that the beamscan axis drawn on the head (labeled "1" and "2"), not the slit axis is what sets the direction of the field you are sampling. Data for both vertical and horizontal spatial modes of the beam are included. There is some ellipticity to the beam that I will have to account for in choosing my (new?) mode matching solution.

 

Error bars are included on the green point, they are just small.

Attachment 1: VertPSLScan.pdf
VertPSLScan.pdf
Attachment 2: HorPSLScan.pdf
HorPSLScan.pdf
  186   Tue Jul 21 18:45:26 2009 ConnorLaserFiberalignment of optics

The 1W NPRO is, for the moment, being used for both the FS and Gyro experiments. Aidan and I began aligning the optics for the FS experiment today.

The beam is first passed through a quarter-waveplate to linearize its polarization. A 50-50 beamsplitter picks off the beam for the Gyro experiment. The transmitted beam (for FS) is then attenuated by a half-waveplate and PBS. We steer the transmitted beam through a lens, a Faraday Isolator, and another lens to collimate it.

A photograph of the setup is attached.

We plan to direct the beam through an AOM and couple it to the fibre so we can characterize the acquired phase noise.

Attachment 1: FSprelimsetup.jpg
FSprelimsetup.jpg
  185   Tue Jul 21 02:09:45 2009 DmassLaserPSLPSL Realigned, now better

Quote:

Koji came down to the dungeon to help me with the "what is the junk coming out of the PSL" problem.

We noticed that the beam right after the PBS looked funky, and the beam before the power amplifier stage inside the PSL looked fine.

We decided to realign the optics inside the PSL. We only touched the two steering mirrors immediately preceeding the power amplifier.

We got a 35% increase in power and the output mode looked cleaner to the eye. (No more solar flares)

We are now happy.

 

 The following image contains approximations of the changes we made to the PSL steering mirrors (~5 degree accuracy)

Attachment 1: PSLRealignment.png
PSLRealignment.png
  184   Mon Jul 20 23:19:49 2009 ranaLaserPSLRazor Blades!
Yeah, real micrometer with a reading is > plain screw.
  183   Sun Jul 19 09:44:09 2009 AidanComputingGeneralAliases added for standard LIGO tools

 

I've updated the /home/controls/.bashrc file to include some aliases to standard ligo tools. These should now be accessible anywhere from the command line.

alias dtt="/bin/diaggui"
alias dmt="/bin/dmtviewer"
alias foton="/bin/foton"
alias dv="dataviewer"
 
StripTool is already accessible from the command line.

  182   Fri Jul 17 15:39:31 2009 DmassLaserPSLRazor Blades!

Quote:

Since the beamscan is in a questionable state of scanny goodness, Koji advised that I do some razor blade occlusion power measurements of the beam and then fit an erf to it to find the waist. I took data with the tinker toys pictured below.

I will compare these results with some beamscans results to verify (hopefully) that the beamscan is outputting useful results, not lies.

 Using the setups in the quoted post, I took manual beamscans 23" in front of the PSL enclosure, before the first steering mirror, and fit P = A erf(B*x + C)

 I measured the relative position of my razorblade with a micrometer and calculated the error from an estimated uncertainty of it's angle. This seemed to agree with repeatability of measurements for a given experimental state. Uncertainty < .002"

I watched the Power meter for ~ 60 seconds for each measurement, it fluctuated around some point and seemed to not be drifting @ DC, so the upper and lower error bars of each point included are the bounds of the fluctuation of the power meter. These were less than +/- 5mW about my points, so a fractional uncertainty of about 2% at my maximum power.

 

I got waists of:

Vertical: 799 +/- 4.5 microns

Horizontol 827 +/- 1.2 microns

Attached plot includes data w/ error and functional form of fit

 

As expected, my Chi^2 is "bad" since I am fitting the input beam to the PMC with a 00 mode description of the waist, which ignores all higher order modal content.

Attachment 1: RazorScans.pdf
RazorScans.pdf
  181   Thu Jul 16 23:21:40 2009 DmassLaserPSLPSL Diode Powers and temperatures

At Anamaria's request, I recorded the diode powers (and temperatures) of the PSL.

~5 min after power up they were:

temp/power

20C  / 29.8W

23.8C / 28.6W

24C / 25.2W

25C / 27.8W

6 hours later they were

20C / 29.0W

23.8C / 28.1W

24C / 24.8W

25C 27.3W

ELOG V3.1.3-